A forex dealer in Bangkok is buying US dollars (USD) in exchange for Thai Baht (THB) at a rate of USD 1 = THB 25.0 and selling Chinese yuan (CNY) at a rate of CNY 1 = THB 3.0 . A Bank of China branch next door is selling USD for CNY at a rate of USD 1 = CNY 8.2.

A speculator plans to earn an arbitrage profit by first going to the dealer, where she will exchange USD to THB and then immediately convert the THB she receives to CNY. After this she will go to the bank and convert the CNY she got from the dealer into USD. If she starts with USD 1,000, how much USD profit will she earn if she does this one time?
Answer #1

CNY 1 = THB 3.0
So, THB 1 = CNY (1/3)

USD 1 = CNY 8.2
So, 1 CNY = USD (1/8.2)

If she starts with USD 1000 then
USD 1 = THB 25
So, USD 1,000 = THB 25*(1000) = THB 25,000

THB 1 = CNY (1/3)
So, THB 25,000 = CNY (25,000/3)

1 CNY = USD (1/8.2)
So, CNY (25,000/3) = USD (25,000/3)*(1/8.2) = USD 25,000/24.6 = USD 1,016.26

So, she started by USD 1000 and ended with USD 1,016.26
So, USD profit = 1016.26 - 1000 = USD 16.26